OpenStudy (anonymous):

f(x) = −x^2 + 3x on [0,3]. What is its maximum value?

OpenStudy (perl):

f ' (x) = -2x + 3 solve f '(x) = 0 -2x + 3 = 0 , x = 3/2 compare: f(0) = 0 f(3/2)= 2.25 f(3) = 0 so the absolute max is 2.25

OpenStudy (perl):

on the domain *